OpenStudy (anonymous):

Part 1: Find the area of a rectangular object which has a length of 4 square root of 2 inches and a width of 6 square root of 6 inches. (3 points) Part 2: Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

i guess you are supposed to multiply to get the area \[4\sqrt{2}\times 6\sqrt{6}=24\sqrt{12}\] and since \(\sqrt{12}=2\sqrt{3}\) you can write the answer as \(48\sqrt{3}\)

OpenStudy (anonymous):

the "complete sentences" are up to you

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

part 1. 1st multiply the 2. 4√2 • 6√6 = 24√12 =48√3 part 2. i multiplied the 2 radicals. the outside numbers multiplies and the inside numbers multiplied. then i simplifies the radical. final answer is 48√3. part 3- a piece of paper could have these measurements